### v2

# I-SHEEP: Self-Alignment of LLM from Scratch through an Iterative Self-Enhancement Paradigm

Abstract
Large Language Models (LLMs) have achieved significant advancements, however, the common learning paradigm treats LLMs as passive information repositories, neglecting their potential for active learning and alignment. Some approaches train LLMs using their own generated synthetic data, exploring the possibility of active alignment. However, there is still a huge gap between these one-time alignment methods and the continuous automatic alignment of humans. In this paper, we introduce I-SHEEP, an Iterative Self-EnHancEmEnt Paradigm. This human-like paradigm enables LLMs to continuously self-align from scratch with nothing. Compared to the one-time alignment method Dromedary (Sun et al. 2023b), which refers to the first iteration in this paper, I-SHEEP can significantly enhance capacities on both Qwen and Llama models. I-SHEEP achieves a maximum relative improvement of 78.2% in the Alpaca Eval, 24.0% in the MT Bench, and an absolute increase of 8.88% in the IFEval accuracy over subsequent iterations in Qwen-1.5 72B model. Additionally, I-SHEEP surpasses the base model in various standard benchmark generation tasks, achieving an average improvement of 24.77% in code generation tasks, 12.04% in TrivialQA, and 20.29% in SQuAD. We also provide new insights based on the experiment results. Our codes, datasets, and models are available at https://github.com/multimodal-art-projection/I-SHEEP.
Introduction
Large Language Models (LLMs) have achieved significant success, yet they remain far from achieving the autonomous self-regulation and coherent self-understanding characteristic of human-like Artificial General Intelligence (AGI). In the pretraining phase, LLMs learn and memorize common knowledge from vast amounts of raw text. During the Supervised Fine-Tuning (SFT) phase, LLMs are trained with question-answer pairs to develop instruction-following capabilities. Both phases currently treat LLMs as passive information repositories, neglecting their potential for active learning and alignment. Furthermore, aligning a base model from scratch requires substantial high-quality data across various tasks and scenarios. This data collection, annotation, and cleaning process is labor-intensive, time-consuming, and must ensure diversity and representativeness. Humans can automatically answer relevant questions after acquiring knowledge (Khader et al. 2016), and there is no need for an explicit process like SFT to teach humans how to answer questions. Therefore, exploring active, automatic, and continuous alignment from scratch, akin to human learning, is a crucial step for LLMs toward achieving AGI.
There are several approaches developed to align base models in low-resource scenarios, aiming to reduce or eliminate reliance on human supervision signals (Wang et al. 2022b; Sun et al. 2023b, a; Xu et al. 2024b). For instance, Self Instruct and Dromedary (Wang et al. 2022b; Sun et al. 2023b) actively improve themselves using self-generated data, while Magpie (Xu et al. 2024b) aligns the base model by prompting more powerful models to generate instruction-output pair data. Although these methods demonstrate some level of proactivity, there remains a significant gap between these one-time alignment processes and the continuous, automatic alignment for various environments seen in human learning.
Educational research suggests that metacognitive self-assessment plays a vital role in continuous alignment, helping students reflect on their knowledge and skills, manage cognitive resources, and improve their performance (Yan et al. 2023). Inspired by this perspective, we explore a new human-like paradigm for LLMs, I-SHEEP, designed to enable LLMs to proactively, automatically, and continuously align from scratch, even with minimal external signal. As shown in Figure 1, I-SHEEP begins with seed data and leverages LLMs powerful understanding and generation capabilities to create additional instruction-output pair data. We then perform self-assessment, allowing the model to monitor and assess its learning process. By filtering out incorrect cognitions and retaining accurate ones, LLMs can self-align by training themselves with these correct cognitions. Through iterative repetition of this process, the model can continuously and autonomously align from scratch, relying solely on its internal knowledge.
The main contributions can be summarized as follows: (1) We propose I-SHEEP, a human-like learning paradigm for LLMs, enabling active, automatic, and continuous self-alignment from scratch using only their internal knowledge. Based on the one-time alignment approach, I-SHEEP explores the incorporation of metacognitive self-assessment to monitor and manage the learning process as well as continuous self-alignment from scratch. (2) We integrate metacognitive self-assessment into the alignment process, allowing the model to monitor and manage its learning process. Experiments explore how varying levels of metacognitive abilities impact the efficiency of the I-SHEEP framework. (3) We find that the self-enhancement potential is closely related to the metacognitive self-assessment level and model size. We conducted experiments to explore the performance of I-SHEEP on different model sizes and varying levels of metacognitive self-assessment.
Related Work
Automatic Data Selection
Zhou et al.; Bai et al. emphasize that dataset quality outweighs quantity during the instruction fine-tuning stage. As a result, some studies on instruction data selection have emerged, focusing on identifying high-quality subsets from candidate datasets(Li et al. 2023a; Du, Zong, and Zhang 2023; Liu et al. 2023; Li et al. 2024; Ge et al. 2024; Xia et al. 2024). These methods aim to improve the model performance, accelerate the training process, and facilitate data-efficient alignment. Li et al. introduce an Instruction-Following Difficulty (IFD) metric and use it to select the top 5% of data for fine-tuning models. The filtering phase in the I-SHEEP framework does not rely on carefully selected metrics, external models or human assistance. Additionally, our method is orthogonal to these selection approaches.
Synthetic Data for Improving Model
Generating synthetic data refers to using the powerful generative capabilities of LLMs to create new data that simulates potential real-world scenarios, reducing the need for costly manual labeling. Some methods use the model’s self-generated data to improve itself (Wang et al. 2022b; Sun et al. 2023b, a; Yehudai et al. 2024). Other methods leverage powerful closed models to generate synthetic data, enhancing the capabilities of open-source models (Taori et al. 2023; Chiang et al. 2023; Xu et al. 2023a; Yu et al. 2023; Wei et al. 2023). In addition to generating complete instruction-output pairs, some methods collect existing raw data and synthesize corresponding questions or answers to create supervised data for improving the model (Huang et al. 2022; Li et al. 2023b; Zheng et al. 2024b; Mitra et al. 2024; Wang et al. 2022a; Asai et al. 2023). Some methods begin with instruction-output pairs, generating feedback or refining answers to improve data quality and enhance the model’s reasoning capabilities.(Lu et al. 2023; Li and He 2024; Gou et al. 2023). The I-SHEEP framework evolves from the aforementioned static, one-time improvement paradigm to a dynamic, continuous self-enhancement process.
Iterative Enhancement for LLMs
There are several approaches to iterative enhancement that rely on the help of strong models or external tools (Chen et al. 2024, 2023; Lu et al. 2023; Gao et al. 2023; Lee et al. 2024). IterAlign (Chen et al. 2024) employs strong models like GPT-4 and Claude2 to detect and correct errors in responses from base LLMs and give the corresponding constitution for improving the safety of LLMs. These methods in iterative enhancement typically depend on strong models or external tools to guarantee ongoing model optimization and avoid model collapse. In addition, some methods explore iterative enhancement in the RLHF phase to continuously align the model with human preference (Yuan et al. 2024; Liu et al. 2024; Pang et al. 2024; Xu et al. 2024a, 2023b; Wu et al. 2024; Wang et al. 2024). These iterative RLHF methods start with the aligned model, while we focus on the base model continuous self-alignment from scratch.
Methodology
Self-Driven Data Synthesis
Self Instruct (Wang et al. 2022b) leverages an off-the-shelf large language model (LLM) for the generation of synthetic data. The approach starts with a small set of 175 prompts, known as the seed task pool, leveraging the model’s powerful understanding and generative capabilities to generate a broader range of prompts and responses. This section elaborates on the Self-Driven Data Synthesis process from two perspectives: Instruction generation and response generation. For ease and consistency in data creation, we utilize a standardized instruction format introduced by Alpaca (Taori et al. 2023), enabling the direct generation of instructions along with their corresponding potential inputs.
Instruction generation. Having some prompts from the seed dataset and the meta-prompt from Alpaca (Taori et al. 2023). The process that model generating new prompt set through In-Context Learning (ICL) can be modeled as:
denotes a new prompt generated by model , represents a subset sampled from the seed dataset for in-context learning (ICL). The symbol stands for the parameter of model .
Response generation. After obtaining the set of prompts , we use the model to generate corresponding responses via a zero-shot approach.
Self-Assessment and Data Filtering
To ensure that the data used for self-enhancement maintains a high-quality standard, a two-stage process comprising self-assessment and data filtering is implemented.
Self-Assessment. We pair the generated prompt set and response set to form the instruction-output pair data . Given the capacity limitations of models, ensuring the quality of synthetic pairs can be challenging, making it essential to assess the quality of the generated data. Manual assessment is often impractical, therefore, we introduce an automated assessment method that relies solely on the model. Specifically, the model autonomously evaluates each generated response for its quality and adherence to the instructions. Each entry is scored based on predefined criteria, which quantitatively reflect the compliance and quality of the response.
Data Filtering. After the self-assessment, the subsequent data filtering phase discards entries that do not meet the specified quality threshold. This step guarantees that only entries of the highest quality are retained in the dataset, thereby enhancing the overall reliability and utility of the generated data. Initially, we apply heuristic rule-based filtering to the generated data during data generation, following the Self-Instruct (Wang et al. 2022b). Additionally, after data generation, we filter the instruction-output pairs based on the assessment scores from the self-assessment phrase. A threshold is applied to filter based on assessment scores, yielding a high-quality dataset .
Input: Initial seed task set , Base model
Hyper-parameter: Iteration steps , Filtering threshold , Data size
Output: Enhanced LLMs , High-quality datasets
Iterative Continuous Model Enhancements
The Iterative Self-Enhancement algorithm aims to incrementally enhance a language model by generating and utilizing high-quality synthetic datasets. As shown in Algorithm 1, starting with an initial model and a small seed task set , the algorithm iterates over a specified number of steps and a filtering threshold . At each iteration , the algorithm performs several functions: it generates a new set of prompts, , using a prompt generation process that leverages the current model and the seed data . It then produces corresponding responses, , forming a raw dataset, = . This dataset undergoes a self-assessment process to evaluate the quality of responses, after which it is filtered using the threshold to retain only high-quality data, resulting in . The model is then trained on to align it closely with the refined data, enhancing its performance iteratively by supervised fine-tuning (SFT) approach. This process continues until it concludes at step , ultimately producing a stronger language model and a refined synthetic dataset .
Experiments
Evaluation
Chat Evaluation
We evaluate the instruction-following ability and response quality of aligned models with three chat benchmarks, AlpacaEval(Dubois et al. 2023), MT-Bench(Zheng et al. 2024a), and IFEval(Zhou et al. 2023), due to their comprehensiveness, fine granularity, and reproducibility. Both AlpacaEval and MT-Bench rely on GPT as an evaluator. IFEval provides four types of accuracy scores: prompt-level strict-accuracy, inst-level strict-accuracy, prompt-level loose-accuracy, and inst-level loose-accuracy.
OpenCompass Evaluation
We use the OpenCompass evaluation platform (Contributors 2023), a comprehensive one-stop platform for LLM evaluation. The evaluation includes standard benchmarks such as BoolQ (Clark et al. 2019), PIQA (Bisk et al. 2019), SIQA (Sap et al. 2019), HellaSwag (Zellers et al. 2019), WinoGrande (Sakaguchi et al. 2019), ARC-c (Clark et al. 2018), OpenBookQA-Fact (Mihaylov et al. 2018), CommonsenseQA (Contributors 2023), and MMLU (Hendrycks et al. 2020). It also includes code generation benchmarks such as HumanEval (Chen et al. 2021) and MBPP (Austin et al. 2021), word knowledge benchmark TriviaQA (Joshi et al. 2017), and reading comprehension benchmark SQuAD2.0 (Rajpurkar, Jia, and Liang 2018). Full results on these benchmarks are available in Appendix CC.
Main Settings
We conduct experiments on the Qwen-1.5 (Team 2024) and Llama-3 (Dubey et al. 2024) models to validate the effectiveness and generalization of I-SHEEP. Additionally, we explore the impact of different model sizes on I-SHEEP by conducting experiments on Qwen-1.5 1.8B, 4B, 7B, 14B, 32B, and 72B models, providing a detailed analysis based on the experimental results. In each iteration, the dataset for training is generated by the model from the last iteration. The case study of the generated data and the overall quality analysis can be found in Appendix BB and Appendix FF, respectively. We utilized LLaMA-Factory (Zheng et al. 2024c) for LoRA fine-tuning, with specific parameters detailed in Appendix EE. Under the configuration of using VLLM for inference (Kwon et al. 2023), the maximum duration of each iteration is about 4 hours on NVIDIA A800-SXM4-80GB8, equivalent to one iteration time for Qwen-1.5 72B.
Self-Assessment and Filter Settings
During the self-assessment phase, we propose three variants, simple standard prompt, combined standard prompt, and ICL prompt, to evaluate data quality. Detailed prompt contents can be found in appendix AA.
In the filtering phase, there are six settings, simple standard prompt based filtering, combined standard prompt based filtering, ICL prompt filtering, PerPLexity (PPL) filtering, density filtering, and the combination of density and PPL filtering. In addition to the first three filtering settings based on scores obtained in the Self-Assessment phase, we also explore data filtering methods that do not rely on external tools or models. For example, PPL filtering uses the PPL value computed by the model itself to evaluate the quality of instruction-output pairs, thereby eliminating low-quality data. We filter out data points with PPL greater than 50. Density filtering extracts vector representations from the model’s final layer and performs K Nearest Neighbors (KNN) clustering, sampling from each cluster to ensure dataset diversity. We set 3000 as the clustering number K. The combination of density and PPL filtering setting first clusters the data and then selects samples with lower PPL values from each cluster, ensuring the filtered dataset’s quality and diversity.
Baseline
We use the base model, Self Instruct (Wang et al. 2022b), and Dromedary (Sun et al. 2023b) as baselines to explore the continuous and automatic enhancement of the human-like framework, I-SHEEP. Self Instruct is a one-time alignment approach where LLMs are trained directly on data they generate, without a self-assessment phase. Similarly, Dromedary is a one-time alignment process where the model generates responses following specific principles, which are then engraved into the model. This approach is similar to the first iteration setting described in this paper.
Iterative Settings and Ablation Settings
Iterative Settings. We investigate the impact of I-SHEEP on efficiency across different iterative self-enhancement settings, including using data generated by the last iteration model to train the base model, using data generated by the last iteration model to train the last iteration model, and using data generated by all previous iterations to train the base model. Additionally, we directly generate 20K and 30K data points for comparative experiments to eliminate the influence of data size in the iterative settings mentioned above. Notably, in the first iteration, all settings are identical, where the base model generates 10k data, filters it, and uses it to fine-tune itself, akin to the Dromedary(Sun et al. 2023b).
| Setting | Chat Benchmark | Standard Benchmark | |||||||||
| Alpaca Eval | MT Bench | IFEval | Code | Knowledge | Reading Comprehension | ||||||
| P-level S-accuracy | I-level S-accuracy | P-level L-accuracy | I-level l-accuracy | Human Eval/Plus | MBPP | Trivia QA | SQuAD 2.0 | ||||
| 1.8B | base | – | – | – | – | – | – | 6.71/6.10 | 16.40 | 31.18 | 30.02 |
| iter1 | 1.51 | 3.76 | 15.53 | 25.30 | 17.74 | 28.06 | 11.59/9.15 | 16.80 | 19.38 | 13.16 | |
| iter2 | 1.54 | 3.53 | 16.27 | 27.10 | 19.22 | 31.41 | 15.24/12.20 | 17.40 | 16.88 | 14.57 | |
| iter3 | 2.30 | 3.16 | 13.68 | 24.46 | 15.34 | 27.22 | 14.02/10.98 | 17.80 | 12.49 | 13.91 | |
| 4B | base | – | – | – | – | – | – | 10.98/8.54 | 28.00 | 40.95 | 27.96 |
| iter1 | 2.61 | 4.97 | 19.41 | 29.98 | 24.03 | 34.77 | 30.49/26.83 | 34.00 | 38.94 | 24.90 | |
| iter2 | 2.96 | 4.79 | 19.78 | 32.61 | 23.84 | 36.81 | 31.10/27.44 | 35.20 | 37.20 | 24.63 | |
| iter3 | 3.78 | 4.99 | 18.85 | 31.41 | 22.18 | 35.37 | 32.93/28.66 | 35.80 | 35.37 | 31.67 | |
| 7B | base | – | – | – | – | – | – | 10.98/8.54 | 36.60 | 51.00 | 33.14 |
| iter1 | 5.19 | 5.08 | 28.47 | 39.93 | 31.05 | 43.41 | 45.73/39.63 | 41.20 | 45.81 | 26.36 | |
| iter2 | 5.37 | 5.13 | 30.13 | 40.89 | 33.09 | 43.88 | 47.56/42.68 | 41.00 | 42.83 | 28.36 | |
| iter3 | 5.22 | 4.97 | 29.21 | 40.29 | 30.68 | 43.05 | 45.12/40.24 | 40.60 | 40.53 | 33.76 | |
| 14B | base | – | – | – | – | – | – | 17.68/15.85 | 41.40 | 57.72 | 20.37 |
| iter1 | 4.77 | 5.68 | 28.84 | 41.13 | 33.46 | 46.40 | 45.73/40.85 | 49.00 | 56.81 | 30.52 | |
| iter2 | 6.27 | 5.97 | 30.87 | 42.93 | 33.46 | 46.40 | 48.78/42.07 | 45.60 | 54.45 | 38.57 | |
| iter3 | 7.30 | 5.48 | 30.13 | 43.05 | 33.27 | 46.04 | 50.00/43.29 | 45.20 | 55.30 | 43.42 | |
| 32B | base | – | – | – | – | – | – | 22.56/21.34 | 47.40 | 65.88 | 29.56 |
| iter1 | 8.27 | 5.56 | 33.46 | 45.32 | 37.52 | 50.12 | 58.54/51.83 | 44.20 | 60.81 | 41.34 | |
| iter2 | 8.26 | 5.68 | 36.04 | 47.60 | 39.56 | 51.92 | 56.71/50.61 | 41.80 | 59.43 | 42.15 | |
| iter3 | 9.30 | 5.69 | 36.41 | 47.96 | 38.82 | 51.56 | 56.71/51.83 | 42.20 | 59.73 | 44.04 | |
| iter4 | 8.64 | 5.62 | 33.83 | 46.88 | 38.45 | 51.56 | 56.10/50.61 | 40.60 | 58.95 | 47.07 | |
| 72B | iter1 | 6.64 5.19 | 6.43 1.54 | 35.67 8.88 | 49.16 6.72 | 40.48 7.02 | 53.96 4.79 | 50.61/45.12 6.10/8.54 | 51.20 4.80 | 60.81 9.62 | 50.68 17.27 |
| iter2 | 9.06 | 7.90 | 37.34 | 51.32 | 40.85 | 54.56 | 56.71/49.39 | 51.80 | 61.55 | 52.27 | |
| iter3 | 10.51 | 7.97 | 41.22 | 54.32 | 44.18 | 57.19 | 56.10/50.61 | 52.60 | 62.00 | 61.42 | |
| iter4 | 11.22 | 5.45 | 42.14 | 54.56 | 46.21 | 58.63 | 51.83/47.56 | 56.00 | 70.43 | 64.55 | |
| iter5 | 11.83 | 5.62 | 44.55 | 55.88 | 47.50 | 58.75 | 56.71/53.66 | 55.60 | 70.11 | 67.95 | |
| iter6 | 11.60 | 5.75 | 42.33 | 53.84 | 45.10 | 56.95 | 51.22/48.17 | 55.20 | 70.01 | 67.82 | |
| Base Model | – | – | – | – | – | – | 21.34/20.12 35.37/33.54 | 50.20 5.80 | 58.07 12.36 | 47.66 20.29 | |
| Self Instruct | 5.26 6.57 | 7.82 0.15 | 33.64 10.91 | 47.60 8.28 | 39.56 7.94 | 53.00 5.75 | 53.05/46.95 3.66/6.71 | 48.40 7.60 | 71.25 -0.82 | 51.90 16.05 | |
Ablation Settings. we adjust high-dimensional variables such as the threshold in the self-assessment phase, data size in the generation phase, and iteration steps in the iterative training phase to validate their impact on I-SHEEP. Furthermore, we conduct ablation experiments with different levels of metacognitive self-assessment, including no self-assessment, assessing only response quality, assessing only instruction-following degree, and assessing both response quality and instruction-following degree.
Results
Main Results
Table 1 shows the experimental performance of various model sizes across different iteration steps. There are some new findings: (1) I-SHEEP exhibits efficacy across various model sizes, with particularly notable improvements in 72B. I-SHEEP achieves a maximum relative improvement of 78.2% in the Alpaca Eval, 24.0% in the MT Bench, and an absolute increase of 8.88% in the IFEval prompt-level strict accuracy over subsequent iterations in Qwen-1.5 72B model. Additionally, I-SHEEP surpasses the base model in various standard benchmark generation tasks, achieving an average improvement of 24.77% in code generation tasks, 12.04% in Trivial QA, and 20.29% in SQuAD. we find that the scores for the second round of dialogues drop significantly after the fourth iteration. This decline is likely due to our generated data consisting solely of single-round dialogues, which do not improve and may even harm the scores for the second round of dialogues. More analysis can be found in the Appendix DD. (2) The potential for improvement varies with different model sizes. The 1.8B, 4B, 7B, and 14B models exhibit improvements over two iterations, 32B and 72B model can improve three and five iterations, respectively, according to the IFEval benchmark.
Iterative Setting Results
| Setting | Chat Benchmark | ||||||
| Alpaca Eval | MT Bench | IFEval | |||||
| P-level S-accuracy | I-level S-accuracy | P-level L-accuracy | I-level L-accuracy | ||||
| iter1(Dromedary) | 6.64 | 6.43 | 35.67 | 49.16 | 40.48 | 53.96 | |
| Direct | 20k | 7.18 | 7.87 | 39.37 | 50.72 | 43.25 | 54.56 |
| 30k | 6.53 | 7.75 | 38.08 | 50.24 | 43.07 | 54.92 | |
| Total_base | iter2 | 7.25 | 7.94 | 39.00 | 50.72 | 45.47 | 56.47 |
| iter3 | 7.51 | 7.94 | 37.52 | 48.32 | 41.59 | 52.76 | |
| One_last | iter2 | 7.76 | 7.76 | 38.45 | 50.48 | 41.96 | 54.92 |
| iter3 | 8.45 | 7.82 | 38.63 | 51.80 | 42.70 | 56.12 | |
| One_base | iter2 | 9.06 | 7.90 | 37.34 | 51.32 | 40.85 | 54.56 |
| iter3 | 10.51 | 7.97 | 41.22 | 54.32 | 44.18 | 57.19 | |
Table 2 presents the chat benchmark performance for the Qwen-1.5 72B model across various iteration settings. More benchmark results are available in Appendix C C. Our findings are as follows: (1) Training the base model with data from the last iteration model is effective for iterative self-enhancement. At the third iteration in the One_base Setting, training the base model with the last iteration data achieves the highest performance on the chat benchmark. The notable performance improvement under this setting suggests that the model has the potential for further enhancement (refer to Table 1 72B results). Therefore, we chose the One_base setting for all subsequent experiments. (2) The data size is not the main factor influencing iterative improvement. Training the base model with the last iteration data at the 3rd iteration outperforms training the base model with a combination of all data from previous iterations.
Threshold Ablation
As shown in Figure 2(a), as the threshold increases, the performance of I-SHEEP at the 3rd iteration shows an upward trend. The threshold 8 is selected to ensure the possibility of further iterative improvement, given the significant performance increase in iteration 2 and iteration 3, and the good performance at iteration 3 with a threshold of 8. Choosing a threshold of 8 is not necessarily the optimal experimental setting, as thresholds of 6, 7, 8, and 9 are all possible.
| Setting | Chat Benchmark | ||||
| IFEval | |||||
| P-level S-accuracy | I-level S-accuracy | P-level L-accuracy | I-level L-accuracy | ||
| Density | iter1 | 34.20 | 46.76 | 39.56 | 51.80 |
| iter2 | 37.34 | 49.76 | 41.22 | 53.72 | |
| iter3 | 37.52 | 49.52 | 39.56 | 51.56 | |
| PPL | iter1 | 36.60 | 49.16 | 41.77 | 54.08 |
| iter2 | 36.04 | 46.64 | 39.92 | 50.84 | |
| iter3 | 33.27 | 45.92 | 36.41 | 49.52 | |
| Density and PPL | iter1 | 37.52 | 49.64 | 42.51 | 54.68 |
| iter2 | 40.48 | 52.16 | 44.73 | 56.24 | |
| iter3 | 38.82 | 50.48 | 41.96 | 53.60 | |
| Simple Standard Prompt | iter1 | 35.30 | 48.20 | 42.33 | 54.68 |
| iter2 | 36.23 | 49.28 | 40.67 | 53.60 | |
| iter3 | 42.14 | 54.08 | 45.10 | 56.83 | |
| Combined Standard Prompt | iter1 | 35.67 | 49.16 | 40.48 | 53.96 |
| iter2 | 37.34 | 51.32 | 40.85 | 54.56 | |
| iter3 | 41.22 | 54.32 | 44.18 | 57.19 | |
| ICL Prompt | iter1 | 38.82 | 49.40 | 43.99 | 55.04 |
| iter2 | 37.34 | 50.84 | 43.25 | 56.47 | |
| iter3 | 41.22 | 53.72 | 43.99 | 36.12 | |
Data Size Ablation
Figure 2(b) shows a stable improvement in the first three iterations across different data sizes (10k, 20k, 30k, 40k, 50k), demonstrating the robustness of the I-SHEEP framework with respect to data size. When the data size is 10k, the model performs well in the 3rd iteration, meanwhile, there are significant improvements between the first iterations. Considering the above factors and resource savings, we chose 10k as the final data size setting.
Metacognitive Self-Assessment Analysis
Self-Assessment Robustness Analysis
Table 3 shows the performance of various self-assessment degrees in the first three iterations. See the Appendix CC for more benchmark results. The following findings can be drawn from the table: (1) Using explicit self-assessment prompt is better than using simple model internal states. On all four IFEval accuracies, the highest values are obtained in the setting where the model is explicitly prompted for self-assessment. (2) The I-SHEEP framework is robust to prompt. Although the criteria differ between simple and combined standard prompt settings, their performance is quite similar. Even without designing a prompt, using just a few examples for ICL can achieve comparable results.
| Setting | IFEval | |||
| P-level S-accuracy | I-level S-accuracy | P-level L-accuracy | I-level L-accuracy | |
| no_prompt_iter1 | 35.67 | 47.60 | 41.04 | 52.88 |
| no_prompt_iter2 | 36.97 | 48.80 | 40.30 | 51.80 |
| no_prompt_iter3 | 37.52 | 48.92 | 39.37 | 50.72 |
| quality_iter1 | 37.34 | 48.20 | 42.51 | 52.64 |
| quality_iter2 | 36.04 | 49.04 | 40.67 | 53.00 |
| quality_iter3 | 37.71 | 51.44 | 41.96 | 54.92 |
| following_iter1 | 35.49 | 47.72 | 38.82 | 51.68 |
| following_iter2 | 40.48 | 52.76 | 43.62 | 56.35 |
| following_iter3 | 39.93 | 51.68 | 43.25 | 55.52 |
| both_iter1 | 35.30 | 48.20 | 42.33 | 54.68 |
| both_iter2 | 36.23 | 49.28 | 40.67 | 53.60 |
| both_iter3 | 41.14 | 54.08 | 45.10 | 56.83 |
Self-Assessment Level Analysis.
As shown in Table 4, we explore the efficiency of I-SHEEP across various self-assessment levels. Our findings include the following key points: (1) The higher the level of self-assessment, the greater the improvement in the efficiency and potential of the I-SHEEP framework. Assessing both quality and instruction-following degree achieves the best performance at 3rd iteration, compared to the other settings. (2) Evaluating the degree of instruction adherence of data pairs is better than only evaluating the quality of output. Compared to the quality experimental group, the following experimental group achieved an overall victory at 2nd iteration on the IFEval benchmark.
Generalization of I-SHEEP
| Setting | IFEval | |||
| P-level S-accuracy | I-level S-accuracy | P-level L-accuracy | I-level L-accuracy | |
| llama3_iter1 | 9.43 | 19.06 | 10.35 | 21.70 |
| llama3_iter2 | 9.61 0.18 | 21.34 2.28 | 11.28 0.93 | 23.74 2.04 |
| llama3_iter3 | 12.38 2.95 | 20.98 1.92 | 14.42 4.07 | 23.86 2.16 |
we conduct experiments on the llama 3 70B model to verify that the I-SHEEP framework is also effective for other models. Table 5 shows that llama 3 is also stably and iteratively enhanced through the I-SHEEP framework. Moreover, the significant improvement between the 2nd iteration and the 3rd iteration indicates that llama3 has the potential for further enhancement.
Conclusion
In this paper, we emphasize and formally introduce a challenging task, continuous self-alignment with nothing, which aims to explore how to achieve and to what extent self-alignment can be realized. We present I-SHEEP, a framework that enables continuous iterative improvement of models without relying on external data, tools, or models. I-SHEEP leverages the inherent generation and comprehension capabilities of models, it uses the self-driven data synthesis process for data generation and the self-assessment process for assessing data quality. Based on these assessment scores, high-quality data is filtered and used to train the model itself. Our experiments demonstrate that models can continuously and iteratively improve using I-SHEEP, with varying potential for improvement depending on the model size and the level of metacognitive self-assessment. Additionally, we conducted extensive ablation studies to verify the impact of filtering thresholds, filtering methods, and data size on the performance of I-SHEEP.
Limitation and Future Work
While the I-SHEEP framework can enhance model performance, the extent of final improvement after the RLHF phase remains uncertain. The complete self-improvement process (SFT+RLHF) needs further investigation, which we leave to future work. Additionally, there are increasing ethical concerns about using synthetic data, as it may intensify biases and harmful content in model responses. Although this paper employs strict filtering for generated data to reduce incorrect cognition, it cannot eliminate them. This issue also persists in models aligned with human-annotated data. Furthermore, the prompt evaluation standards we propose can be replaced with safety-related standards to enhance model safety, similar to Dromedary (Sun et al. 2023b).
Appendix A Appendix A – Self-Assessment Prompt Content
In this section, we provide a detailed overview of the different prompts used during the self-assessment phase. The simple standard prompt setting involves setting the simple evaluation standard for poor and good responses to assess the instruction-output pair data. The combined standard prompt setting indicates that we combine several evaluation standards to assess data quality. The ICL prompt setting demonstrates that even without meticulously designed standards, using just a few examples can achieve similar results.
Appendix B Appendix B – Case Study
| Poor quality and not instruction-following |
| "instruction": "Imagine you’re talking to a doctor. You can ask them to show you a radiograph of your knees. Tell them, ’Show me my knees!’", |
| "input": "", |
[…]
| "quality_score": 1, |
| "following_score": 1 |
| "instruction": "Please fill in the blank with a good response.", |
| "input": "_______ is better than ______.", |
| "output": "One", |
| "quality_score": 2, |
| "following_score": 1 |
| Good quality but not instruction-following |
| "instruction": "What are the main strengths and weaknesses of this technology?", |
| "input": "Technology: Self-driving cars", |
[…]
| "quality_score": 10, |
| "following_score": 2 |
| "instruction": "Create a new formula for generating a unique ID.", |
| "input": "", |
[…]
| "quality_score": 8, |
| "following_score": 5 |
| Poor quality but instruction-following |
| "instruction": "Put parentheses around the following expression so that it evaluates to 15.", |
| "input": "5 + 3 * 2", |
| "output": "(15.5 + 3) * 2", |
| "quality_score": 2, |
| "following_score": 9 |
| "instruction": "Remove the spaces from this text.", |
| "input": "Here is a text with spaces.", |
[…]
| "quality_score": 1, |
| "following_score": 10 |
| "instruction": "You will be given a task to generate a creative response to the given prompt. A creative response is something that is interesting and novel.", |
| "input": "A group of friends want to have a fun night in.", |
| "output": "They could play board games or watch a movie.", |
| "quality_score": 4, |
| "following_score": 8 |
| Good quality and instruction-following |
| "instruction": "Explain the following joke.", |
| "input": "Why do programmers prefer dark mode? Because light attracts bugs.", |
| "output": "The joke is a play on words. ’Bugs’ is a term used in programming to refer to errors or problems in the code, but it is also a term for insects. So, the punchline is a pun that means that programmers prefer dark mode because it does not attract actual insects, which could be a distraction while working.", |
| "quality_score": 10, |
| "following_score": 10 |
| "instruction": "Explain the grammar knowledge for the given sentence.", |
| "input": "An ancient vase stands on a brown table.", |
[…]
| "quality_score": 10, |
| "following_score": 10 |
Appendix C Appendix C – More Benchmark Results Evaluated by Opencompass
In this section, we present more benchmark results evaluated using the Opencompass platform. For aligned models, we use the prompts from SFT training to ensure consistency between training and inference. The prompts used are as follows:
| dataset | version | metric | mode | Qwen base model | 1st iteration | 2nd one_base | 3rd one_base |
| Standard Benchmarks | |||||||
| BoolQ | 314797 | accuracy | ppl | 89.45 | 89.24 | 89.30 | 89.54 |
| piqa | 0cfff2 | accuracy | ppl | 83.35 | 83.24 | 83.24 | 83.08 |
| siqa | e8d8c5 | accuracy | ppl | 77.89 | 78.35 | 78.40 | 78.51 |
| GPQA_diamond | 152005 | accuracy | gen | 25.25 | 27.78 | 26.77 | 27.78 |
| hellaswag | a6e128 | accuracy | ppl | 83.45 | 83.39 | 83.46 | 83.46 |
| winogrande | 55a66e | accuracy | ppl | 75.30 | 75.14 | 74.82 | 74.66 |
| ARC-e | 2ef631 | accuracy | ppl | 96.12 | 96.12 | 96.30 | 96.12 |
| ARC-c | 2ef631 | accuracy | ppl | 91.86 | 92.20 | 91.53 | 90.85 |
| openbookqa_fact | 6aac9e | accuracy | ppl | 94.40 | 94.80 | 95.00 | 95.60 |
| commonsense_qa | e51e32 | accuracy | ppl | 77.23 | 77.56 | 77.97 | 77.89 |
| mmlu | - | naive_average | ppl | 77.02 | 76.85 | 76.95 | 77.03 |
| Code Generation | |||||||
| openai_humaneval | 812847 | pass@1 | gen | 21.34 | 50.61 | 56.71 | 56.10 |
| mbpp | d1bbee | score | gen | 50.20 | 51.20 | 51.80 | 52.60 |
| World Knowledge | |||||||
| nq | 632c4e | score | gen | 19.11 | 26.54 | 27.31 | 28.14 |
| triviaqa | f9d2af | score | gen | 58.07 | 60.81 | 61.55 | 62.00 |
| Reading Comprehension | |||||||
| squad2.0 | 817436 | score | gen | 47.66 | 50.68 | 52.27 | 61.42 |
| dataset | version | metric | mode | Qwen base model | 1st iteration | 2nd one_last | 3rd one_last |
| Standard Benchmarks | |||||||
| BoolQ | 314797 | accuracy | ppl | 89.45 | 89.24 | 89.30 | 89.20 |
| piqa | 0cfff2 | accuracy | ppl | 83.35 | 83.24 | 83.13 | 82.92 |
| siqa | e8d8c5 | accuracy | ppl | 77.89 | 78.35 | 78.25 | 78.56 |
| GPQA_diamond | 152005 | accuracy | gen | 25.25 | 27.78 | 27.27 | 28.28 |
| hellaswag | a6e128 | accuracy | ppl | 83.45 | 83.39 | 83.39 | 83.37 |
| winogrande | 55a66e | accuracy | ppl | 75.30 | 75.14 | 75.37 | 75.14 |
| ARC-e | 2ef631 | accuracy | ppl | 96.12 | 96.12 | 96.30 | 96.30 |
| ARC-c | 2ef631 | accuracy | ppl | 91.86 | 92.20 | 92.20 | 91.86 |
| openbookqa_fact | 6aac9e | accuracy | ppl | 94.40 | 94.80 | 95.00 | 95.60 |
| commonsense_qa | e51e32 | accuracy | ppl | 77.23 | 77.56 | 78.05 | 77.89 |
| mmlu | - | naive_average | ppl | 77.02 | 76.85 | 76.86 | 76.95 |
| Code Generation | |||||||
| openai_humaneval | 812847 | pass@1 | gen | 21.34 | 50.61 | 56.71 | 56.10 |
| mbpp | d1bbee | score | gen | 50.20 | 51.20 | 51.80 | 52.60 |
| World Knowledge | |||||||
| nq | 632c4e | score | gen | 19.11 | 26.54 | 27.31 | 28.14 |
| triviaqa | f9d2af | score | gen | 58.07 | 60.81 | 61.55 | 62.00 |
| Reading Comprehension | |||||||
| squad2.0 | 817436 | score | gen | 47.66 | 50.68 | 52.27 | 61.42 |
| dataset | version | metric | mode | Qwen base model | 1st iteration | 2nd total_base | 3rd total_base |
| Standard Benchmarks | |||||||
| BoolQ | 314797 | accuracy | ppl | 89.45 | 89.24 | 89.17 | 89.27 |
| piqa | 0cfff2 | accuracy | ppl | 83.35 | 83.24 | 83.19 | 83.19 |
| siqa | e8d8c5 | accuracy | ppl | 77.89 | 78.35 | 78.20 | 78.25 |
| GPQA_diamond | 152005 | accuracy | gen | 25.25 | 27.78 | 27.27 | 26.26 |
| hellaswag | a6e128 | accuracy | ppl | 83.45 | 83.39 | 83.43 | 83.47 |
| winogrande | 55a66e | accuracy | ppl | 75.30 | 75.14 | 75.14 | 75.22 |
| ARC-e | 2ef631 | accuracy | ppl | 96.12 | 96.12 | 96.30 | 96.30 |
| ARC-c | 2ef631 | accuracy | ppl | 91.86 | 92.20 | 91.86 | 91.86 |
| openbookqa_fact | 6aac9e | accuracy | ppl | 94.40 | 94.80 | 95.20 | 95.00 |
| commonsense_qa | e51e32 | accuracy | ppl | 77.23 | 77.56 | 77.81 | 77.81 |
| mmlu | - | naive_average | ppl | 77.02 | 76.85 | 76.90 | 76.92 |
| Code Generation | |||||||
| openai_humaneval | 812847 | pass@1 | gen | 21.34 | 50.61 | 56.71 | 56.10 |
| mbpp | d1bbee | score | gen | 50.20 | 51.20 | 51.80 | 52.60 |
| World Knowledge | |||||||
| nq | 632c4e | score | gen | 19.11 | 26.54 | 27.31 | 28.14 |
| triviaqa | f9d2af | score | gen | 58.07 | 60.81 | 61.55 | 62.00 |
| Reading Comprehension | |||||||
| squad2.0 | 817436 | score | gen | 47.66 | 50.68 | 52.27 | 61.42 |
| dataset | version | metric | mode | Qwen base model | 1st iteration | direct 20K | direct 30K |
| Standard Benchmarks | |||||||
| BoolQ | 314797 | accuracy | ppl | 89.45 | 89.24 | 89.20 | 89.54 |
| piqa | 0cfff2 | accuracy | ppl | 83.35 | 83.24 | 83.35 | 83.24 |
| siqa | e8d8c5 | accuracy | ppl | 77.89 | 78.35 | 77.79 | 78.15 |
| GPQA_diamond | 152005 | accuracy | gen | 25.25 | 27.78 | 26.77 | 25.76 |
| hellaswag | a6e128 | accuracy | ppl | 83.45 | 83.39 | 83.43 | 83.44 |
| winogrande | 55a66e | accuracy | ppl | 75.30 | 75.14 | 75.37 | 75.14 |
| ARC-e | 2ef631 | accuracy | ppl | 96.12 | 96.12 | 96.47 | 96.30 |
| ARC-c | 2ef631 | accuracy | ppl | 91.86 | 92.20 | 90.85 | 91.53 |
| openbookqa_fact | 6aac9e | accuracy | ppl | 94.40 | 94.80 | 95.00 | 94.80 |
| commonsense_qa | e51e32 | accuracy | ppl | 77.23 | 77.56 | 77.72 | 77.40 |
| mmlu | - | naive_average | ppl | 77.02 | 76.85 | 76.96 | 76.97 |
| Code Generation | |||||||
| openai_humaneval | 812847 | pass@1 | gen | 21.34 | 50.61 | 56.71 | 56.10 |
| mbpp | d1bbee | score | gen | 50.20 | 51.20 | 51.80 | 52.60 |
| World Knowledge | |||||||
| nq | 632c4e | score | gen | 19.11 | 26.54 | 27.31 | 28.14 |
| triviaqa | f9d2af | score | gen | 58.07 | 60.81 | 61.55 | 62.00 |
| Reading Comprehension | |||||||
| squad2.0 | 817436 | score | gen | 47.66 | 50.68 | 52.27 | 61.42 |
| Setting | Chat Benchmark | Standard Benchmark | |||||||
| IFEval | Code | World Knowledge | Reading Comprehension | ||||||
| Prompt-level Strict-accuracy | Inst-level Strict-accuracy | Prompt-level Loose-accuracy | Inst-level loose-accuracy | Human Eval/Plus | MBPP | Trivia QA | SQuAD 2.0 | ||
| Density | iter1 | 34.20 | 46.76 | 39.56 | 51.80 | 53.66/46.34 | 50.60 | 70.95 | 53.50 |
| iter2 | 37.34 | 49.76 | 41.22 | 53.72 | 51.83/44.51 | 53.40 | 70.78 | 60.58 | |
| iter3 | 37.52 | 49.52 | 39.56 | 51.56 | 54.88/47.56 | 55.20 | 69.97 | 59.54 | |
| PPL | iter1 | 36.60 | 49.16 | 41.77 | 54.08 | 52.44/46.95 | 50.00 | 71.34 | 50.40 |
| iter2 | 36.04 | 46.64 | 39.92 | 50.84 | 56.71/50.00 | 52.20 | 70.27 | 48.11 | |
| iter3 | 33.27 | 45.92 | 36.41 | 49.52 | 55.49/50.61 | 53.20 | 70.37 | 41.82 | |
| Density and PPL | iter1 | 37.52 | 49.64 | 42.51 | 54.68 | 52.44/46.95 | 50.60 | 71.29 | 57.08 |
| iter2 | 40.48 | 52.16 | 44.73 | 56.24 | 55.49/48.17 | 54.40 | 70.87 | 62.06 | |
| iter3 | 38.82 | 50.48 | 41.96 | 53.60 | 58.54/53.05 | 55.40 | 70.40 | 63.51 | |
| Simple Standard Prompt | iter1 | 35.30 | 48.20 | 42.33 | 54.68 | 53.66/46.34 | 51.20 | 71.39 | 51.51 |
| iter2 | 36.23 | 49.28 | 40.67 | 53.60 | 56.71/50.00 | 55.60 | 71.17 | 57.64 | |
| iter3 | 42.14 | 54.08 | 45.10 | 56.83 | 59.76/53.05 | 57.60 | 70.40 | 63.47 | |
| Combined Standard Prompt | iter1 | 35.67 | 49.16 | 40.48 | 53.96 | 50.61/45.12 | 51.20 | 60.81 | 50.68 |
| iter2 | 37.34 | 51.32 | 40.85 | 54.56 | 56.71/49.39 | 51.80 | 61.55 | 52.27 | |
| iter3 | 41.22 | 54.32 | 44.18 | 57.19 | 56.10/50.61 | 52.60 | 62.00 | 61.42 | |
| ICL Prompt | iter1 | 38.82 | 49.40 | 43.99 | 55.04 | 54.27/47.56 | 53.40 | 71.45 | 58.62 |
| iter2 | 37.34 | 50.84 | 43.25 | 56.47 | 59.76/53.05 | 54.60 | 71.49 | 57.91 | |
| iter3 | 41.22 | 53.72 | 43.99 | 36.12 | 59.15/52.44 | 55.40 | 69.88 | 58.91 | |
Appendix D Appendix D – MT-Bench
| single turn score | coding | extraction | humanities | math | reasoning | role play | stem | writing | total average | ||
| iter1 | 1st turn | 7.76 | 5.50 | 6.35 | 9.65 | 5.85 | 7.60 | 7.55 | 9.55 | 10.00 | 6.43 |
| 2nd turn | 5.11 | 2.30 | 4.80 | 7.70 | 3.60 | 5.30 | 7.90 | 4.30 | 5.00 | ||
| iter2 | 1st turn | 8.23 | 7.10 | 7.90 | 9.60 | 6.10 | 7.40 | 8.30 | 9.90 | 9.55 | 7.90 |
| 2nd turn | 7.57 | 5.05 | 8.30 | 9.70 | 4.90 | 8.00 | 9.00 | 7.80 | 7.80 | ||
| iter3 | 1st turn | 8.34 | 6.50 | 7.80 | 9.65 | 7.00 | 7.20 | 8.80 | 10.00 | 9.80 | 7.97 |
| 2nd turn | 7.60 | 5.80 | 7.60 | 9.40 | 5.00 | 7.50 | 9.30 | 8.50 | 7.70 | ||
| iter4 | 1st turn | 7.43 | 5.00 | 7.70 | 9.70 | 4.95 | 5.80 | 7.10 | 9.40 | 9.75 | 5.45 |
| 2nd turn | 3.48 | 2.40 | 2.40 | 5.40 | 2.30 | 2.40 | 5.80 | 4.50 | 2.60 | ||
| iter5 | 1st turn | 7.49 | 5.30 | 7.70 | 9.50 | 4.90 | 5.60 | 7.80 | 9.40 | 9.70 | 5.62 |
| 2nd turn | 3.76 | 3.60 | 2.50 | 5.20 | 1.50 | 3.60 | 4.80 | 4.40 | 4.50 | ||
| iter6 | 1st turn | 7.74 | 5.10 | 7.00 | 9.45 | 7.80 | 5.60 | 7.80 | 9.50 | 9.70 | 5.75 |
| 2nd turn | 3.73 | 3.00 | 2.50 | 7.10 | 2.20 | 3.60 | 5.30 | 3.11 | 3.00 |
Appendix E Appendix E –Lora Hyperparameters and LLaMA Factory Template
We present the hyperparameters used for LoRA training and the templates used for SFT in the LLama-Factory framework as follows:
Appendix F Appendix F – Data quality analysis across different iterations